OpenStudy (anonymous):

Factor Completely: x2 − 49 (x − 7)(x − 7) (x + 7)(x − 7) (x + 7)(x + 7) Prime

Parth (parthkohli):

Difference of squares - \(( a+ b)( a- b) = a^2 - b^2\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 - 7^2 = ( x +7)(\cdots)}\)

OpenStudy (anonymous):

would my answer be the second one?

Parth (parthkohli):

Yea